OpenStudy (anonymous):

What are the zeroes of f(x) = x2 - 6x + 8? (4 points) x = -4, 2 x = -4, -2 x = 4, 2 x = 4, -2

OpenStudy (anonymous):

you can use quadratic formula or you can find the factors of the coefficients of the leading term and the last term.

OpenStudy (ybarrap):

$$ f(x)=x^2-6x+8=(x-2)(x-4) $$ When is $$ f(x)=0 $$ That is, when is x-2 =0 and x-4 = 0?

OpenStudy (anonymous):

the substitute those values into the equation, until you get zero

OpenStudy (anonymous):

B?

OpenStudy (bee_see):